OpenStudy (anonymous):

find the general solution of the differential equation dy dx = (2 cos x − 3) e(−3x+2 sin x)/6 y5/6

OpenStudy (anonymous):

In the implicit form

OpenStudy (anonymous):

this is the third time you've posted this lol

OpenStudy (anonymous):

i'll do a detailed solution

OpenStudy (anonymous):

Yes. I have some problem with it still. but nearly there.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

so this is the equation right? \[\frac{dy}{dx} = (2cosx -3)e^{\frac{2sinx - 3x}{6}} y^{\frac{5}{8}}\]

OpenStudy (anonymous):

we are going to separate variables we want anything with a y in it on the LHS

OpenStudy (anonymous):

y5/6

OpenStudy (anonymous):

so divide by y^(5/6) \[y^{\frac{-5}{6}}\frac{dy}{dx} = (2cosx -3)e^{\frac{2sinx - 3x}{6}}\]

OpenStudy (anonymous):

y is y^5/6

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now we are going to integrate both sides with respect to x \[\int\limits{y^{\frac{-5}{6}}\frac{dy}{dx}}\text{ }dx = \int\limits{(2cosx -3)e^{\frac{2sinx - 3x}{6}}} \text{ } dx\]

OpenStudy (anonymous):

on the left hand side, we have the chain rule going on, because we have dy/dx multiplied by a function of y in fact this chain rule effect allows us to rewrite that left integral such that we are integrating with respect to y: \[\int\limits{y^{\frac{-5}{6}}} dy = \int\limits{(2cosx -3)e^{\frac{2sinx - 3x}{6}}} \text{ } dx\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

lets evaluate the left integral right now: \[6y^{\frac{1}{6}} = \int\limits{(2cosx -3)e^{\frac{2sinx - 3x}{6}}} \text{ } dx\] notice i haven't added an arbitrary constant just yet. i could have, but when we evaluate the second integral we will get one, and so its easier just to add it in later \[ 6y^{\frac{1}{6}} = \int\limits{(2cosx -3)e^{\frac{2sinx - 3x}{6}}} \text{ } dx\] next we will be evaluating the right integral.

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

\[I =\int\limits{(2cosx -3)e^{\frac{2sinx - 3x}{6}}} \text{ } dx\] we can do this by substitution: let \[u = 2sinx - 3x\] so \[\frac{du}{dx} = 2cosx - 3\]\[dx = \frac{du}{2cosx - 3}\] \[I = \int\limits{(2cosx -3)e^{\frac{u}{6}}}dx = \int\limits{(2cosx -3)e^{\frac{u}{6}}}\frac{du}{2cosx - 3}\] \[I = \int\limits{e^\frac{u}{6}}\text{ }du\]

OpenStudy (anonymous):

following?

OpenStudy (anonymous):

Ok. So (2cosx -3) gets cancelled out?

OpenStudy (anonymous):

What is the method call by which you are doing this? composite rule or substitution method?

OpenStudy (anonymous):

it's integration by substitution. if you find it difficult then all it is is practise so we get: \[6y^{\frac{1}{6}} = \int\limits{e^\frac{u}{6}}\text{ }du = \frac{1}{6} e^{\frac{u}{6}} + C\]

OpenStudy (anonymous):

all we need to do now is solve for y: tip: multiplying or adding to C by a constant does not really affect it, as we just get another arbitrary constant. \[6y^{\frac{1}{6}} = \frac{1}{6} e^{\frac{u}{6}} + C\]\[y^{\frac{1}{6}} = \frac{1}{36} e^{\frac{u}{6}} + C\]\[y = (\frac{1}{36} e^{\frac{u}{6}} + C)^6\]

OpenStudy (anonymous):

Is this the implicit form or explicit form? It seems explicit.

OpenStudy (anonymous):

I have tried to use this to resolve the following initial values but it does not seem right. That is why I posted it again to see if it was correct.

OpenStudy (anonymous):

Find the particular solution of the differential equation for which y = 1 when x = 0, and then give this particular solution in explicit form.

OpenStudy (anonymous):

hmm it does look a bit explicit. remember to sub back in u: \[36y = (e^{\frac{-3x + 2sinx}{6}}+C)^6\] i guess to make it implicit you could square both sides.. if you want another one to practise the techniques here: \[\frac{dy}{dx} = 2xy^2e^{(x^2)}\]

OpenStudy (anonymous):

i just thought, squaring both sides probably not a good idea, it doubles up the curve

OpenStudy (anonymous):

With Y = 1 and x= 0 . I am trying to find the value for C. But not convinced by what I get.